Consider the following results of a simple regression model of dollar price of unleaded gas (dependent...

  1. Consider the following results of a simple regression model of dollar price of unleaded gas (dependent variable) and dollar price of crude oil (independent variable) and answer the following questions:

Coefficient

t-statistics

Intercept

0.56

28.27

Crude Oil

0.0457

73.34

R-square = 0.87

a. What is the interpretation of coefficient for crude oil?

Every 1% increase in crude price will cause unleaded gas price to increase by 4.57 cents.

Every $1 increase in crude price will cause unleaded gas price to increase by 4.57 cents.

Every 1% increase in crude price will cause unleaded gas price to increase by 4.57%.

Every $1 increase in crude price will cause unleaded gas price to increase by 4.57 %.


b. What will be expected change in the price of unleaded gas if the crude oil price is expected to fall by 3%. The current price of crude oil and unleaded gas are $74 and $3.10 respectively?

(3.27%)

(2.25%)

(3.85%)

(2.89%)

c. What will be forecasted price of unleaded gas when the crude oil price is $90?

$4.44

$4.67

$4.22

$3.99
Answer #1

Every $1 increase in crude price will cause unleaded gas price to increase by 4.57 cents.

=(-74*3%*0.0457)/3.10=-3.273%

=0.56+0.0457*90=4.673